OpenStudy (meilendurcer):

Mai-li has 6 1/3 yd of material. Her new skirt will take 2 1/3 yd. How much material will she have left after the skirt is made? A. 8 yd B. 8 2/3 yd C. 4 yd D. 12 2/3

OpenStudy (anonymous):

okay so you are going to subtract like this

OpenStudy (anonymous):

can you try this one

OpenStudy (meilendurcer):

sure

hero (hero):

\[\space \space 6\frac{1}{3}\]\[-2\frac{1}{3}\]

OpenStudy (anonymous):

let me know if you stuck this one is real easy just ignore the fraction so you have left 6-2

hero (hero):

Me personally, though, I'd rather show it like this because it is more obvious: \[(6-2) + (\frac{1}{3} - \frac{1}{3})\]

OpenStudy (anonymous):

I learned the other way along time ago and I have had a hard time doing it any other way

OpenStudy (meilendurcer):

@Hero 26?

hero (hero):

What you learned is called the vertical method. Mine is the horizontal method

OpenStudy (anonymous):

no what is 6-2

hero (hero):

2 different forms of the same thing @moser90

OpenStudy (meilendurcer):

i hate math, i don't even know it well

OpenStudy (anonymous):

what is 6-2

OpenStudy (anonymous):

that is the answer to the problem